# Lean-STaR: Learning to Interleave Thinking and Proving

### v3

@@ Abstract
Traditional language model-based theorem proving assumes that by training on a sufficient amount of formal proof data, a model will learn to prove theorems. Our key observation is that a wealth of informal information that is not present in formal proofs can be useful for learning to prove theorems. For instance, humans think through steps of a proof, but this thought process is not visible in the resulting code. We present Lean-STaR, a framework for training language models to produce informal thoughts prior to each step of a proof, thereby boosting the model’s theorem-proving capabilities. Lean-STaR uses retrospective ground-truth tactics to generate synthetic thoughts for training the language model. At inference time, the trained model directly generates the thoughts prior to the prediction of the tactics in each proof step. Building on the self-taught reasoner framework, we then apply expert iteration to further fine-tune the model on the correct proofs it samples and verifies using the Lean solver. Lean-STaR achieves state-of-the-art results on the miniF2F-test benchmark within the Lean theorem proving environment, significantly outperforming base models ($\bm{43.4\%\rightarrow 46.3\%,}$ Pass@64). We also analyze the impact of the augmented thoughts on various aspects of the theorem proving process, providing insights into their effectiveness.
@@ 1 Introduction
Theorem proving is a fundamental aspect of mathematics, and mathematical reasoning is an important part of artificial intelligence [28, 55]. Formalized mathematics in particular provides a challenging testbed for assessing mathematical reasoning capabilities. Since theorems and proofs in this setting can be represented in the form of checkable source code, it is easy to evaluate proofs of arbitrary complexity [15]. Automated theorem proving, if successful, can also help discover unknown errors in previous proofs [^1] For example, Terence Tao found a non-trivial error while using Lean to formalize a project [38]., and make it easier to guarantee that new proofs are correct. More broadly, formal mathematics coupled with powerful automation may unlock new forms of education and collaboration, mathematical insights, and applications to verifying critical software [4, 16, 10, 31].
@@
Recently, language models have shown promising progress in formal theorem proving [33, 35, 47, 19, 25, 50, 26]. Existing approaches typically train a model solely based on the proofs in a formal language (code) such as Lean [15], Isabelle [29], or Coq [13]. Our key observation is that such approaches ignore a wealth of informal information that may be useful for learning to prove theorems [45, 46]. For instance, the underlying thought process prior to each step of a proof is not present in formal source code. Based on this insight, we propose to train a language model that can produce a natural language chain-of-thought (“thought”) prior to each step (“tactic”) of a formal proof.
@@
We introduce Lean-STaR, a framework for learning to interleave informal thoughts with steps of formal proving. Building on the Self-Taught Reasoner (STaR) framework [52], we enable language models to interleave step-by-step rationales (i.e., thoughts) [30, 43] with formal proving in a two-stage process. In an initial phase, we prompt a sufficiently capable language model, such as GPT-4 [1], and generate retrospective thoughts based on a dataset of human-written proofs, such as Mathlib, the largest collection of human-written proofs in Lean [27]. Subsequently, we fine-tune a thought-augmented tactic predictor [8, 7, 18, 14] that, given a Lean state, can generate a thought and predict the subsequent tactic. In a second phase, we optimize this thought-augmented tactic predictor with the expert iteration algorithm [3, 36], using multi-step success rate in theorem proving as the reward.
@@
Our work presents a new link between informal and formal mathematics, complementary to prior explorations that translate standalone mathematical statements [37, 42, 48] or translate informal proofs into formal proofs [2, 23, 5, 56, 22]. Lean-STaR generates natural language thoughts specifically for each proof step, improving formal proving capabilities by interleaving natural and formal languages.
@@
We instantiate Lean-STaR by generating roughly 50,000 thought-augmented examples from Lean’s Mathlib [27], then synthesize an additional 50k examples through two iterations of expert iteration. To the best of our knowledge, this yields the first thought-augmented dataset for theorem proving. After fine-tuning an InternLM2-7b base model [51] on our thought-augmented data, our final Lean-STaR model can solve $\mathbf{34.8\%}$ (pass@32) or $\mathbf{36.1\%}$ (pass@64) of the problems on miniF2F-test [54]. Using stronger base model InternLM2-7b-plus, Lean-STaR can achieve $\mathbf{45.4\%}$ (pass@32), significantly surpassing the previous state-of-the-art results of $43.4\%$ (pass@32). In summary, Lean-STaR offers a framework for teaching language models to interleave informal thoughts with formal verification, advancing the capabilities of language models in automated theorem proving.
@@ 2 Related Work
Previous work on learning-based theorem proving typically follows the GPT-f framework [33], which trains a language model on (proof state, next-tactic) pairs, then proves theorems by using the model within a best-first tree search. Subsequent work has explored several directions, including data augmentation [20], novel proof search methods [25, 41], further training through curriculum learning [34], retrieval augmentation [50], or practical tools [44]. Others use prompted models to generate tactics [6, 39], or fine-tune models to generate a full proof [17]. A second auto-formalization [48] thread incorporates informal mathematics into formal theorem proving. Draft-Sketch-Prove [23] shows that language models have some ability to use informal proofs to improve a model’s formal proving abilities, by drafting an informal proof, translating into a formal proof sketch, then completing the proof with tools like Sledgehammer [8]. Draft-Sketch-Prove and related methods [40, 53, 57] are limited to the Isabelle prover, since they use powerful automatic proving tools like Sledgehammer. Lean lacks these tools, so generating the entire proof at once would be more unlikely in Lean. We focus on Lean, and train language models to generate a thought and predict the subsequent tactic in each proof step. To the best of our knowledge, we are the first to introduce thought-augmented reasoning in automatic theorem proving.
@@
Recently, many works demonstrated that letting language models reason before an answer can improve their performance on tasks including math, science, and code [30, 43, 12]. Although the corresponding techniques (e.g., Scratchpad and Chain-of-Thought) have proven to be effective, they require either extensive annotated training examples or exposure to numerous similar examples during pre-training [9]. The scarcity of natural language reasoning in formal theorem proving, coupled with the impracticality of manually annotating rationales for formal mathematics, thus presents a challenge. We propose a new Lean-STaR framework for synthesizing training examples by taking advantage of the correctness signal from the formal system.
@@
Recently, several works suggest that language models may be taught to reason via synthetic data that they generate themselves, akin to a reinforcement learning method that improves a policy through self-play. Polu et al. [34] showed that a simple RL algorithm, expert iteration, paired with curriculum learning can improve a formal theorem proving model. Self-Taught Reasoner (STaR) [52] showed that we can iteratively fine-tune the language model on the correct (reasoning, answer) pairs generated by itself to gradually improve performance. Singh et al. [36] proposed ReST-EM, which filters data generated by language model with a binary feedback signal rather than using fully manually annotated data (similar to expert iteration in [34]). Our work builds on these ideas, providing the first study of bootstrapped thought-augmented proving.
@@ 3.1 Preliminaries
Interactive Theorem Provers (ITPs) are typically used for step-by-step automatic theorem proving in formal mathematics. At each step, we can provide the ITP with a high-level “tactic” to simplify the current goal state (e.g., the initial goal theorems to be proven) into subgoals. These subgoals will form new states, and proving all the subgoals results in a complete proof of the given theorem. We use Lean [15], a popular interactive theorem prover. An example formal proof in Lean and its explanation are shown in Appendix C.
### State
a b n : $\mathbb{N}$
$\vdash$ ¬ (n $\neq$ 0 $\land$ a $\neq$ 0) $\leftrightarrow$ n = 0 $\lor$ a = 0
### Tactic
“‘lean4
simp only [not_and_or, ne_eq, not_not]
### Reasoning
Figure 3: Examples of (input, output) pairs of retrospective rationale generation with GPT-4. The full prompt is given in Appendix F. See §3.2.2 for the detailed explanation.
@@ 3.2.1 Direct Tactic Prediction
In the typical setting [33], proving a theorem consists of providing a proof state $s$ to the language model and then generating a tactic from the language model $M$, i.e., $\pi_{M}(a|s)$. The language model can be fine-tuned for this task using a dataset of (proof state, next-tactic) pairs from successful proof trajectories, i.e. $D=\{(s^{i},a^{i}):i=1,\cdots,M\}$, where final states have a reward of 1. We refer to a language model fine-tuned on such a dataset as a supervised fine-tuning (SFT) model.
@@ 3.2.2 Thought-augmented Tactic Prediction
Existing approaches typically train only on formal states and tactics [33]. We hypothesize that incorporating a latent thought can improve a model’s ability to predict the next tactic. Formally, we introduce a hidden “thought” variable $t_{i}$ prior to each tactic, and then extend the model to the form $\pi_{M}(a_{i},t_{i}|s_{i})=\pi_{M}(a_{i}|t_{i},s_{i})\pi_{M}(t_{i}|s_{i})$. In thought-augmented tactic prediction, the distribution over the next tactic can then be expressed as:
@@
The key challenge is obtaining (state, thought, tactic) pairs for training a model. To this end, we introduce retrospective rationale generation. Our motivating observation is that the distribution of natural language thoughts in theorem-proving $\pi_{M}(t_{i}|s_{i})$ is scarce in the pre-training corpus of large language models. In turn, we find that even the most powerful GPT-4 model does not perform well in generating the correct rationale through few-shot prompting [9]. To develop a language model capable of generating thoughts and tactics $a_{i},t_{i}|s_{i}$, we need an entirely new dataset $D_{T}=\{(s^{i},t^{i},a^{i}):i=1,\cdots,N\}$. However, in Lean, we only have a dataset of $D_{S}=\{(s^{i},a^{i}):i=1,\cdots,N\}$ where $(s^{i},a^{i})$ is one step in some successful proof trajectories. Given a powerful large language model $G$, which we refer to as the oracle model [^2] For instance, in our experiments we use the best available large language model, GPT-4., we give the oracle model the ground-truth tactic $a_{i}$ and let the oracle model produce the thought $t_{i}$ given the current state $s_{i}$ and ground-truth tactic $a_{i}$. This helps improve the pass rate and produce thought-augmented data more efficiently. Our few-shot prompt is provided in Appendix F. The design principle of the prompt is to prevent the oracle model from generating hindsight-like thoughts.
@@ 3.2.3 Bootstrapping Thought-augmented Theorem Proving
This method can be seen as an offline RL method [36] in the theorem proving MDP. In this MDP, the cumulative reward $R\left((s_{0},t_{0},a_{0}),(s_{1},t_{1},a_{1}),\cdots,(s_{n},t_{n},a_{n})\right)=1$ if and only if the proof trajectory is successful. The total expected reward is
@@
and Lean-STaR’s expert iteration can be seen as optimizing this reward [36].
@@ 3.3 Evaluation
The most popular method to evaluate the theorem proving ability of a language model $M$ is to use best-first search like GPT-f [33, 50, 6, 44]. In best-first search, we keep all unexpanded states $s_{i}$. Each time, we expand the “best” state $s_{i}$ and use the language model to sample $S$ next tactics $a_{i,1\cdots S}$ for the current state $s_{i}$. For each legal tactic $a_{i,j}$, a new state can be obtained by applying tactic $a_{i,j}$ on state $s_{i}$. Following standard practice [33, 50, 44], we assume the state with maximum log-probabilities is the “best”s. Specifically, we select state $s_{i}$ with maximum $\sum\limits_{j=0}^{i-1}\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),\cdots,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average log probability of each generated token. We expand up to $N$ states and we get a successful proof search when we reach any proof state with no goals. Then, we can attempt the search $K$ times to obtain a pass rate $pass@K$. However, we found that the best-first search method performs poorly with the Lean-CoT and Lean-STaR models. We attribute this to using average log probabilities, which may not be a reliable quality indicator when the thought sequence $t_{j}$ is generated.
@@ 4 Experiments
We instantiate Lean-STaR using the best available open language model pre-trained on the Lean corpus (InternLM2-Math-base-7b [51]), and follow standard practice in using Lean’s Mathlib as the underlying training set (via the Lean Dojo dataset [50]). We generate an initial set of thoughts for Mathlib using GPT-4, perform two rounds of expert iteration, then evaluate the model on miniF2F [54], the de-facto standard benchmark for evaluating language-model based theorem provers. Our experimental results show that both retrospective rationale generation and expert iteration significantly improve the theorem-proving capabilities of language models in this setting. We describe our setup and findings in detail below.
Table 1: Pass rates on the minif2f-test dataset with Lean. This table shows the pass rates of previous works and our work. $S$ is the number of tactics attempted at each expanded node (assumed to be $1$ in sampling) and $K$ is the total number of search or sampling attempts per problem. In sampling we use temperature 0.7, and in search we use beam search when generating the next tactic. Note that we sample $32$ examples twice when $K=64$ in sampling.
Approach Decoding $N$ $K$ $S$ Pass rate
GPT-3.5 [1] (Few-Shot) Sampling 50 1 1 $2.8\%$
GPT-4 [1] (Few-Shot) Sampling 50 1 1 $11.9\%$
Transformer [34] (w/o RL) Search 512 1 8 $24.6\%$
Llemma-34b [6] (Few-Shot) Search 50 1 32 $25.8\%$
Llemma-7b [6] (Few-Shot) Search 50 1 32 $26.2\%$
ReProver [50] Search 50 1 64 $26.5\%$
Transformer [34] (w/ RL) Search 512 1 8 $29.6\%$
InternLM2-34b [51] (Few-Shot) Search 50 1 32 $29.5\%$
COPRA (with GPT-4) [39] Customized - 60 1 $29.9\%$
COPRA (with GPT-4) [39] Customized - 100 1 $30.7\%$
InternLM2-7b [51] (Few-Shot) Sampling 50 32 1 $28.7\%$
InternLM2-7b [51] (Few-Shot) Search 50 1 32 $30.3\%$
SFT (InternLM2-7b) Sampling 50 32 1 $29.5\%$
@@
Lean-CoT (InternLM2-7b) Sampling 50 32 1 $32.8\%$
Lean-STaR (Iter-1) (InternLM2-7b) Sampling 50 32 1 $34.0\%$
Lean-STaR (Iter-2) (InternLM2-7b) Sampling 50 32 1 $\mathbf{34.8\%}$

### v4

@@ Abstract
Traditional language model-based theorem proving assumes that by training on a sufficient amount of formal proof data, a model will learn to prove theorems. Our key observation is that a wealth of informal information that is not present in formal proofs can be useful for learning to prove theorems. For instance, humans think through steps of a proof, but this thought process is not visible in the resulting code. We present Lean-STaR, a framework for training language models to produce informal thoughts prior to each step of a proof, thereby boosting the model’s theorem-proving capabilities. Lean-STaR uses retrospective ground-truth tactics to generate synthetic thoughts for training the language model. At inference time, the trained model directly generates the thoughts prior to the prediction of the tactics in each proof step. Building on the self-taught reasoner framework, we then apply expert iteration to further fine-tune the model on the correct proofs it samples and verifies using the Lean solver. Lean-STaR significantly outperforming base models ($\bm{43.4\%\rightarrow 46.3\%,}$ Pass@64). We also analyze the impact of the augmented thoughts on various aspects of the theorem proving process, providing insights into their effectiveness.
@@ 1 Introduction
Theorem proving is a fundamental aspect of mathematics, and mathematical reasoning is an important part of artificial intelligence (Newell & Simon, 1956; Zhou, 2023). Formalized mathematics in particular provides a challenging testbed for assessing mathematical reasoning capabilities. Since theorems and proofs in this setting can be represented in the form of checkable source code, it is easy to evaluate proofs of arbitrary complexity (De Moura et al., 2015). Automated theorem proving, if successful, can also help discover unknown errors in previous proofs [^1] For example, Terence Tao found a non-trivial error while using Lean to formalize a project (Tao, 2023)., and make it easier to guarantee that new proofs are correct. More broadly, formal mathematics coupled with powerful automation may unlock new forms of education and collaboration, mathematical insights, and applications to verifying critical software (Avigad, 2023; First, 2023; Buzzard, 2024; of Sciences, 2023).
@@
Recently, language models have shown promising progress in formal theorem proving (Polu & Sutskever, 2020; Rabe et al., 2020; Wu et al., 2021; Han et al., 2021; Lample et al., 2022; Yang et al., 2023; Li et al., 2024). Existing approaches typically train a model solely based on the proofs in a formal language (code) such as Lean (De Moura et al., 2015), Isabelle (Nipkow et al., 2002), or Coq (Coq, 1996). Our key observation is that such approaches ignore a wealth of informal information that may be useful for learning to prove theorems (Welleck et al., 2021; 2022). For instance, the underlying thought process prior to each step of a proof is not present in formal source code. Based on this insight, we propose to train a language model that can produce a natural language chain-of-thought (“thought”) prior to each step (“tactic”) of a formal proof.
@@
We introduce Lean-STaR, a framework for learning to interleave informal thoughts with steps of formal proving. Building on the Self-Taught Reasoner (STaR) framework (Zelikman et al., 2022), we enable language models to interleave step-by-step rationales (i.e., thoughts) (Nye et al., 2021; Wei et al., 2022) with formal proving in a two-stage process. In an initial phase, we prompt a sufficiently capable language model, such as GPT-4 (Achiam et al., 2023), and generate retrospective thoughts based on a dataset of human-written proofs, such as Mathlib, the largest collection of human-written proofs in Lean (mathlib Community, 2020). Subsequently, we fine-tune a thought-augmented tactic predictor (Bohme & Nipkow, 2010; Blanchette et al., 2016; Gloeckle et al., 2023; Czajka & Kaliszyk, 2018) that, given a Lean state, can generate a thought and predict the subsequent tactic. In a second phase, we optimize this thought-augmented tactic predictor with the expert iteration algorithm (Anthony et al., 2017; Singh et al., 2023), using multi-step success rate in theorem proving as the reward.
@@
Our work presents a new link between informal and formal mathematics, complementary to prior explorations that translate standalone mathematical statements (Szegedy, 2020; Wang et al., 2020; Wu et al., 2022) or translate informal proofs into formal proofs (Agrawal et al., 2022; Jiang et al., 2022; Azerbayev et al., 2023a; Zhou et al., 2024a; Huang et al., 2024). Lean-STaR generates natural language thoughts specifically for each proof step, improving formal proving capabilities by interleaving natural and formal languages.
@@
We instantiate Lean-STaR by generating roughly 50,000 thought-augmented examples from Lean’s Mathlib (mathlib Community, 2020), then synthesize an additional 50k examples through two iterations of expert iteration. To the best of our knowledge, this yields the first thought-augmented dataset for theorem proving. After fine-tuning an InternLM2-7b base model (Ying et al., 2024) on our thought-augmented data, our final Lean-STaR model can solve $\mathbf{34.8\%}$ (pass@32) or $\mathbf{36.1\%}$ (pass@64) of the problems on miniF2F-test (Zheng et al., 2021). Using stronger base model InternLM2-7b-plus, Lean-STaR can achieve $\mathbf{45.4\%}$ (pass@32), significantly surpassing the previous results of $43.4\%$ (pass@32). In summary, Lean-STaR offers a framework for teaching language models to interleave informal thoughts with formal verification, advancing the capabilities of language models in automated theorem proving.
@@ 2 Related Work
Previous work on learning-based theorem proving typically follows the GPT-f framework (Polu & Sutskever, 2020), which trains a language model on (proof state, next-tactic) pairs, then proves theorems by using the model within a best-first tree search. Subsequent work has explored several directions, including data augmentation (Han et al., 2022), novel proof search methods (Lample et al., 2022; Wang et al., 2023b), further training through curriculum learning (Polu et al., 2022), retrieval augmentation (Yang et al., 2023), or practical tools (Welleck & Saha, 2023). Others use prompted models to generate tactics (Azerbayev et al., 2023b; Thakur et al., 2023), or fine-tune models to generate a full proof (First et al., 2023). A second auto-formalization (Wu et al., 2022) thread incorporates informal mathematics into formal theorem proving. Draft-Sketch-Prove (Jiang et al., 2022) shows that language models have some ability to use informal proofs to improve a model’s formal proving abilities, by drafting an informal proof, translating into a formal proof sketch, then completing the proof with tools like Sledgehammer (Bohme & Nipkow, 2010). Draft-Sketch-Prove and related methods (Wang et al., 2023a; Zhao et al., 2024; Zhou et al., 2024b) are limited to the Isabelle prover, since they use powerful automatic proving tools like Sledgehammer. Lean lacks these tools, so generating the entire proof at once would be more unlikely in Lean. We focus on Lean, and train language models to generate a thought and predict the subsequent tactic in each proof step. To the best of our knowledge, we are the first to introduce thought-augmented reasoning in automatic theorem proving.
@@
Recently, many works demonstrated that letting language models reason before an answer can improve their performance on tasks including math, science, and code (Nye et al., 2021; Wei et al., 2022; Chen et al., 2022). Although the corresponding techniques (e.g., Scratchpad and Chain-of-Thought) have proven to be effective, they require either extensive annotated training examples or exposure to numerous similar examples during pre-training (Brown et al., 2020). The scarcity of natural language reasoning in formal theorem proving, coupled with the impracticality of manually annotating rationales for formal mathematics, thus presents a challenge. We propose a new Lean-STaR framework for synthesizing training examples by taking advantage of the correctness signal from the formal system.
@@
Recently, several works suggest that language models may be taught to reason via synthetic data that they generate themselves, akin to a reinforcement learning method that improves a policy through self-play. Polu et al. (2022) showed that a simple RL algorithm, expert iteration, paired with curriculum learning can improve a formal theorem proving model. Self-Taught Reasoner (STaR) (Zelikman et al., 2022) showed that we can iteratively fine-tune the language model on the correct (reasoning, answer) pairs generated by itself to gradually improve performance. Singh et al. (2023) proposed ReST-EM, which filters data generated by language model with a binary feedback signal rather than using fully manually annotated data (similar to expert iteration in (Polu et al., 2022)). Our work builds on these ideas, providing the first study of bootstrapped thought-augmented proving.
@@ 3.1 Preliminaries
Interactive Theorem Provers (ITPs) are typically used for step-by-step automatic theorem proving in formal mathematics. At each step, we can provide the ITP with a high-level “tactic” to simplify the current goal state (e.g., the initial goal theorems to be proven) into subgoals. These subgoals will form new states, and proving all the subgoals results in a complete proof of the given theorem. We use Lean (De Moura et al., 2015), a popular interactive theorem prover. An example formal proof in Lean and its explanation are shown in Appendix D.
@@ 3.2.1 Direct Tactic Prediction
In the typical setting (Polu & Sutskever, 2020), proving a theorem consists of providing a proof state $s$ to the language model and then generating a tactic from the language model $M$, i.e., $\pi_{M}(a|s)$. The language model can be fine-tuned for this task using a dataset of (proof state, next-tactic) pairs from successful proof trajectories, i.e. $D=\{(s^{i},a^{i}):i=1,\cdots,M\}$, where final states have a reward of 1. We refer to a language model fine-tuned on such a dataset as a supervised fine-tuning (SFT) model.
@@ 3.2.2 Thought-augmented Tactic Prediction
Existing approaches typically train only on formal states and tactics (Polu & Sutskever, 2020). We hypothesize that incorporating a latent thought can improve a model’s ability to predict the next tactic. Formally, we introduce a hidden “thought” variable $t_{i}$ prior to each tactic, and then extend the model to the form $\pi_{M}(a_{i},t_{i}|s_{i})=\pi_{M}(a_{i}|t_{i},s_{i})\pi_{M}(t_{i}|s_{i})$. In thought-augmented tactic prediction, the distribution over the next tactic can then be expressed as:
@@
The key challenge is obtaining (state, thought, tactic) pairs for training a model. To this end, we introduce retrospective rationale generation. Our motivating observation is that the distribution of natural language thoughts in theorem-proving $\pi_{M}(t_{i}|s_{i})$ is scarce in the pre-training corpus of large language models. In turn, we find that even the most powerful GPT-4 model does not perform well in generating the correct rationale through few-shot prompting (Brown et al., 2020). To develop a language model capable of generating thoughts and tactics $a_{i},t_{i}|s_{i}$, we need an entirely new dataset $D_{T}=\{(s^{i},t^{i},a^{i}):i=1,\cdots,N\}$. However, in Lean, we only have a dataset of $D_{S}=\{(s^{i},a^{i}):i=1,\cdots,N\}$ where $(s^{i},a^{i})$ is one step in some successful proof trajectories. Given a powerful large language model $G$, which we refer to as the oracle model [^2] For instance, in our experiments we use the best available large language model, GPT-4., we give the oracle model the ground-truth tactic $a_{i}$ and let the oracle model produce the thought $t_{i}$ given the current state $s_{i}$ and ground-truth tactic $a_{i}$. This helps improve the pass rate and produce thought-augmented data more efficiently. Our few-shot prompt is provided in Appendix G. The design principle of the prompt is to prevent the oracle model from generating hindsight-like thoughts.
@@ 3.2.3 Bootstrapping Thought-augmented Theorem Proving
This method can be seen as an offline RL method (Singh et al., 2023) in the theorem proving MDP. In this MDP, the cumulative reward $R\left((s_{0},t_{0},a_{0}),(s_{1},t_{1},a_{1}),\cdots,(s_{n},t_{n},a_{n})\right)=1$ if and only if the proof trajectory is successful. The total expected reward is
@@
and Lean-STaR’s expert iteration can be seen as optimizing this reward (Singh et al., 2023).
@@ 3.3 Evaluation
The most popular method to evaluate the theorem proving ability of a language model $M$ is to use best-first search like GPT-f (Polu & Sutskever, 2020; Yang et al., 2023; Azerbayev et al., 2023b; Welleck & Saha, 2023). In best-first search, we keep all unexpanded states $s_{i}$. Each time, we expand the “best” state $s_{i}$ and use the language model to sample $S$ next tactics $a_{i,1\cdots S}$ for the current state $s_{i}$. For each legal tactic $a_{i,j}$, a new state can be obtained by applying tactic $a_{i,j}$ on state $s_{i}$. Following standard practice (Polu & Sutskever, 2020; Yang et al., 2023; Welleck & Saha, 2023), we assume the state with maximum negative log-probabilities is the “best”s. Specifically, we select state $s_{i}$ with maximum $\sum\limits_{j=0}^{i-1}-\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),\cdots,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average log probability of each generated token. We expand up to $N$ states and we get a successful proof search when we reach any proof state with no goals. Then, we can attempt the search $K$ times to obtain a pass rate $pass@K$. However, we found that the best-first search method performed poorly in the Lean-CoT and Lean-STaR models, as detailed in the Appendix E. We attribute this to using average log probabilities, which may not be a reliable quality indicator when the thought sequence $t_{j}$ is generated.
@@ 4 Experiments
We instantiate Lean-STaR using the best available open language model pre-trained on the Lean corpus (InternLM2-Math-base-7b (Ying et al., 2024)), and follow standard practice in using Lean’s Mathlib as the underlying training set (via the Lean Dojo dataset (Yang et al., 2023)). We generate an initial set of thoughts for Mathlib using GPT-4, perform two rounds of expert iteration, then evaluate the model on miniF2F (Zheng et al., 2021) and leandojo (Yang et al., 2023), the de-facto standard benchmark for evaluating language-model based theorem provers. Our experimental results show that both retrospective rationale generation and expert iteration significantly improve the theorem-proving capabilities of language models in this setting. We describe our setup and findings in detail below.
Table 1: Pass rates on the minif2f-test and Leandojo dataset with Lean. This table shows the pass rates of previous works and our work. $S$ is the number of tactics attempted at each expanded node (assumed to be $1$ in sampling) and $K$ is the total number of search or sampling attempts per problem. In sampling we use temperature 0.7, and in search we use beam search when generating the next tactic. We use a random subset of Leandojo4-v9-test (novel premises) with a size of 320 as test set of leandojo. Note that we sample $32$ examples twice when $K=64$ in sampling.
Approach Decoding $N$ $K$ $S$ minif2f leandojo
GPT-3.5 Achiam et al. (2023) (Few-Shot) Sampling 50 1 1 $2.8\%$
GPT-4 Achiam et al. (2023) (Few-Shot) Sampling 50 1 1 $11.9\%$
Transformer Polu et al. (2022) (w/o RL) Search 512 1 8 $24.6\%$
Llemma-34b Azerbayev et al. (2023b) Search 50 1 32 $25.8\%$
Llemma-7b Azerbayev et al. (2023b) Search 50 1 32 $26.2\%$
ReProver Yang et al. (2023) Search 50 1 64 $26.5\%$
Transformer Polu et al. (2022) (w/ RL) Search 512 1 8 $29.6\%$
InternLM2-34b Ying et al. (2024) Search 50 1 32 $29.5\%$
COPRA (with GPT-4) Thakur et al. (2023) Customized - 60 1 $29.9\%$
COPRA (with GPT-4) Thakur et al. (2023) Customized - 100 1 $30.7\%$
InternLM2-7b Ying et al. (2024) Sampling 50 32 1 $28.7\%$ $29.7\%$
InternLM2-7b Ying et al. (2024) Search 50 1 32 $30.3\%$
SFT (InternLM2-7b) Sampling 50 32 1 $29.5\%$ $30.6\%$
@@
Lean-CoT (InternLM2-7b) Sampling 50 32 1 $32.8\%$ $35.6\%$
Lean-STaR (Iter-1) (InternLM2-7b) Sampling 50 32 1 $34.0\%$ $38.4\%$
Lean-STaR (Iter-2) (InternLM2-7b) Sampling 50 32 1 $\mathbf{34.8\%}$ $\mathbf{39.4\%}$